OpenStudy (howard-wolowitz):

METAL for YOUR help :

OpenStudy (howard-wolowitz):

Can we do the first one:

OpenStudy (howard-wolowitz):

what are you thinking dude

OpenStudy (campbell_st):

well in the point you have y = 0 when x = -3.... so y muse be either greater than or equal to zero or less than or equal to... so that eliminates choices 1 and 4 so you have the middle 2 to select from. so substitute x = -3 and y = 0 so is 0 < 6 option 2 or 0 > 6 option 3 you decide

OpenStudy (campbell_st):

the next question is interesting sin < 0 is 3rd of 4th quadrants csc or 1/sin > 0 is 1st of 2nd quadrants which seems to be a conridiction to me.

OpenStudy (campbell_st):

look at 24 24 = 6 x 4 48 = 6 x 8 4 = 2^2 and 8 = 2^3 so it looks like y = 6(2^x)

OpenStudy (howard-wolowitz):

so you think it could be int two quadrants

OpenStudy (howard-wolowitz):

I dont play around little gurl

OpenStudy (anonymous):

I am not a little girl

OpenStudy (kitten_is_back):

http://prntscr.com/77cyhl

OpenStudy (howard-wolowitz):

you think a

OpenStudy (kitten_is_back):

http://prntscr.com/77czah